OpenStudy (anonymous):

How many license plates can be made which have the form ##LLL? 1,757,600 175,600 3,840,500 2,947,000

OpenStudy (chrisasl):

With this plate: 99AAA being a legal (acceptable) one there are 9*9*26*26*26 = 1423656

OpenStudy (anonymous):

I don't want to know how u even computed that... I don't know how to solve this.

OpenStudy (inkyvoyd):

Ok.

OpenStudy (inkyvoyd):

Three numbers. Three letters.

OpenStudy (chrisasl):

There are 9 numbers { 1,2,3,4,5,6,7,8,9 } so 9 choices.. So 9 * 9.. And there are 26 letters in the alphabet..so 26 choices for each letter. 26 * 26 * 26 choices for the 3 letters. All in all there are 9*9*26*26*26 = 1423656

OpenStudy (anonymous):

my god my brain hurts. Somone please just tell me how to solve this.

OpenStudy (chrisasl):

Why do you even not want to know how I solved it? Can't you understand it the way I explained it?

OpenStudy (anonymous):

Not in th slightest.

OpenStudy (anonymous):

100 * 26³ = ..

OpenStudy (chrisasl):

@Chlorophyll, why the 100? @BazBendell: Which part you don't understand? Did you got the numbers?

OpenStudy (anonymous):

I dont understand any of it...

OpenStudy (anonymous):

Chloro's answer gave 1,757,600

OpenStudy (anonymous):

Ten digits 0 - 9 -> 10 * 10 = 100

OpenStudy (chrisasl):

Correct! there's another digit too :P the "0"..I am glad the solution was correct